# Ultrafast Heating Induced Suppression of d-band Dominance in the Electronic Excitation Spectrum of Cuprum

### v1

@@ 3.2 Plasmon dispersion and conditions in interstitial regions
Analyzing the DSF, we have found that the collective plasmon oscillations in X-ray driven copper overcome the dominance of the $d$-band excitations, eventually overtaking them with respect to the spectral weight. In addition, the plasmon position starts to exhibit a substantial dispersion with respect to the wavenumber for sufficiently high temperatures, which is in contrast to the ground state plasmon. To examine the character of the plasmon dispersion, we show the dependence of the plasmon energy (frequency) on the wavenumber, $\omega(q)$, in Fig. 6 for different temperatures. We consider $q<1$-1, where the plasmon peak can be clearly identified at all considered temperatures, and independent of the crystallographic direction. The uncertainty in the plasmon position is evaluated by looking at the onset of a broadened peak. At $T=2~{}{\rm eV}$, we find that $\omega(q)$ is qualitatively similar to the results for $T=0.025~{}{\rm eV}$. For $T\geq 4~{}{\rm eV}$, the plasmon dispersion starts to follow the familiar quadratic dependence on $q$ that is well-known from the free electron gas model 70. To further quantify this trend, we fit the LR-TDDFT results using the Bohm-Gross type dispersion relation 71, 72,
@@
Figure 6: Dependence of the plasmon frequency on the wavenumber along different crystallographic directions at $T=0.025~{}{\rm eV}$, $T=2~{}{\rm eV}$, $T=4~{}{\rm eV}$, $T=6~{}{\rm eV}$, and $T=8~{}{\rm eV}$. Solid (dashed) lines show Bohm-Gross type quadratic dispersion fits [cf. Eq. (6)] for the direction [111] ([100]) at $T=4~{}{\rm eV}$, $T=6~{}{\rm eV}$, and $T=8~{}{\rm eV}$. The horizontal solid grey line at $10.864~{}{\rm eV}$ indicates the plasmon energy computed using the free electron gas model, i.e., assigning the $3d$ shell as a core state 32.
@@
We note that the parameter $\widetilde{r}_{s}$ also characterizes the coupling strength or, equivalently, the degree of non-ideality of the free electrons 34, 73.
An additional, related effective parameter of interest is given by the corresponding degeneracy temperature $\widetilde{\Theta}=k_{B}T/E_{F}(\widetilde{n})$, where $E_{F}(\widetilde{n})=(3\pi^{2}\widetilde{n})^{1/3}$ denotes the Fermi energy of a free electron gas of density $\widetilde{n}$ 43. Finally, we consider the effective ionic charge $\widetilde{Z}$, which is a key ingredient to equation-of-state tables 9, 10. The effective charge is computed using $\widetilde{Z}=\left(\widetilde{r}_{s}(T=0)/\widetilde{r}_{s}(T)\right)^{3}$, which follows from the condition $\widetilde{Z}n_{i}=\widetilde{n}$, with $n_{i}={\rm const}$ being the number density of ions, and setting $Z=1$ for copper at $T=0$ since we have one valence electron in the 4s state.
@@
Upon increasing the temperature, we find a monotonic increase in the density of effectively free electrons, leading to a corresponding decrease of $\widetilde{r}_{s}$ and an increase in $\widetilde{Z}$. The observed monotonic increase of $\widetilde{\Theta}$ is less trivial. On the one hand, we have $\widetilde{\Theta}\sim\widetilde{r}_{s}^{2}$ 73, which, by itself, would indicate a decrease of $\widetilde{\Theta}$ with $T$. However, this effect is overridden by the relation $\widetilde{\Theta}\sim T$ in practice.
Table 1 clearly shows that the effectively free electrons in the interstitial region of isochorically heated copper are in the WDM regime 33. In nature, WDM occurs in a variety of astrophysical objects such as giant planet interiors 3, brown dwarfs 37, and white dwarf atmospheres 36. In the laboratory, these conditions are encountered on the compression path of a fuel capsule and its ablator in inertial confinement fusion experiments 4, 39, 6, and, in addition, they are relevant for material science, synthesis, and discovery 74, 75. Despite its fundamental importance for a gamut of applications, the rigorous theoretical description of WDM remains challenging: it must cover the complex interplay between effects such as partial ionization, Coulomb coupling, and quantum degeneracy and diffraction, which is notoriously difficult in practice 33, 34, 35. Our new simulation results thus imply that future XRTS experiments with isochorically heated copper constitute a suitable and highly controlled testbed for the benchmarking of theoretical methods and simulations.
@@ 4 Conclusions
Motivated by experimental capabilities at modern FEL facilities, we have performed a detailed study of the changes in the electronic DSF in fcc copper due to isochoric heating. Our new LR-TDDFT simulations show that the heating induces a prominent plasmon feature that eventually becomes dominant over the $d$-band-signal. This is in marked contrast to the ground state, where the plasmon is strongly damped by the presence of $d$-band excitations, and where it does not exhibit a meaningful dispersion with respect to the wavenumber $q$. Indeed, we have shown that at $T\geq 4~{}{\rm eV}$, the plasmon dispersion follows the familiar Bohm-Gross type relation, and the plasma frequency $\omega_{p}$ substantially increases with the temperature. This has been explained by the accumulation of effectively free electrons in the interstitial region between the ions. This, in turn, is a consequence of the availability of the electrons in the $d$-orbitals for filling states in the quasicontinuum when the temperature is increased. Interestingly, the reported behavior of the plasmon in copper is fundamentally different from isochorically heated aluminium, where LR-TDDFT simulations have shown that heating to a few electronvolts (at $T\lesssim 7~{}{\rm eV}$) causes a red shift of $\omega_{p}$ 26.
@@
In this regard, we propose to use future XRTS experiments with isochorically heated copper as a controlled testbed for the rigorous assessment of different theoretical models and simulation tools for the description of WDM. Ideally, one might infer the temperature of the heated sample based on the model-free imaginary-time thermometry approach 11, 12 as a first step. Indeed, even the inference of comparably moderate temperatures of $T\sim 1-10\,$eV is expected to be feasible using the new high-resolution set-up that has recently been demonstrated at the European XFEL in Germany 25. Second, we propose to carry out measurements at multiple scattering angles (i.e., multiple wavenumbers $q$) to observe the plasmon dispersion $\omega(q)$ and, in this way, to infer the effective charge state $\widetilde{Z}$ and the effective density parameter $\widetilde{r}_{s}$. Since the ambient density is a-priori known, this will give one full access to the most relevant parameters of the system. This is an important advantage over shock experiments, or XRTS measurements with backlighter sources 40, where one or multiple of these parameters can only be inferred on the basis of the theoretical models which we aim to test in the first place.
Such a hypothetical XRTS dataset can then be used to benchmark the zoo of available theoretical methods such as the widely used effective chemical models 19, 76, 77. A particularly interesting question is the assessment of XC functionals in KS-DFT simulations, including the resolution of explicitly thermal XC effects 78, 79, 80, 81, 82, 83, 84. Moreover, frequency-resolved inelastic X-ray scattering data for a set of finite wavenumbers $q$ will be ideally suited to gauge the accuracy of different XC-kernels in LR-TDDFT calculations, including a rigorous assessment of the popular adiabatic approximation 41, 45, 85.

### v2

@@ 3.2 Plasmon dispersion and conditions in interstitial regions
Analyzing the DSF, we have found that the collective plasmon oscillations in X-ray driven copper overcome the dominance of the $d$-band excitations, eventually overtaking them with respect to the spectral weight. In addition, the plasmon position starts to exhibit a substantial dispersion with respect to the wavenumber for sufficiently high temperatures, which is in contrast to the ground state plasmon. To examine the character of the plasmon dispersion, we show the dependence of the plasmon energy (frequency) on the wavenumber, $\omega(q)$, in Fig. 6 for different temperatures. We consider $q<1$-1, where the plasmon peak can be clearly identified at all considered temperatures, and independent of the crystallographic direction. The uncertainty in the plasmon position is evaluated by looking at the onset of a broadened peak. At $T=2~{}{\rm eV}$, we find that $\omega(q)$ is qualitatively similar to the results for $T=0.025~{}{\rm eV}$. For $T\geq 4~{}{\rm eV}$, the plasmon dispersion starts to follow the familiar quadratic dependence on $q$ that is well-known from the free electron gas model 78. To further quantify this trend, we fit the LR-TDDFT results using the Bohm-Gross type dispersion relation 79, 80,
@@
Figure 6: Dependence of the plasmon frequency on the wavenumber along different crystallographic directions at $T=0.025~{}{\rm eV}$, $T=2~{}{\rm eV}$, $T=4~{}{\rm eV}$, $T=6~{}{\rm eV}$, and $T=8~{}{\rm eV}$. Solid (dashed) lines show Bohm-Gross type quadratic dispersion fits [cf. Eq. (6)] for the direction [111] ([100]) at $T=4~{}{\rm eV}$, $T=6~{}{\rm eV}$, and $T=8~{}{\rm eV}$. In Bohm-Gross dispersion (6) we used for [111] ([100]) $\alpha=1.66\,\omega_{p}^{2}$ ($\alpha=1.7\,\omega_{p}^{2}$) at $T=4~{}{\rm eV}$, $\alpha=1.24\,\omega_{p}^{2}$ ($\alpha=1.3\,\omega_{p}^{2}$) at $T=6~{}{\rm eV}$, and $\alpha=1.11\,\omega_{p}^{2}$ ( $\alpha=1.25\,\omega_{p}^{2}$) at $T=8~{}{\rm eV}$. The horizontal solid grey line at $10.864~{}{\rm eV}$ indicates the plasmon energy computed using the free electron gas model, i.e., assigning the $3d$ shell as a core state [e.g. see Ref. 31].
@@
We note that the parameter $\widetilde{r}_{s}$ also characterizes the coupling strength or, equivalently, the degree of non-ideality of the free electrons 33, 81.
An additional, related effective parameter of interest is given by the corresponding degeneracy temperature $\widetilde{\Theta}=k_{B}T/E_{F}(\widetilde{n})$, where $E_{F}(\widetilde{n})=(3\pi^{2}\widetilde{n})^{1/3}$ denotes the Fermi energy of a free electron gas of density $\widetilde{n}$ 42. Finally, we consider the effective ionic charge $\widetilde{Z}$, which is a key ingredient to equation-of-state tables 9, 10. The effective charge is computed using $\widetilde{Z}=\left(\widetilde{r}_{s}(T=0)/\widetilde{r}_{s}(T)\right)^{3}$, which follows from the condition $\widetilde{Z}n_{i}=\widetilde{n}$, with $n_{i}={\rm const}$ being the number density of ions, and setting $Z=1$ for copper at $T=0$ since we have one valence electron in the 4s state.
@@
Upon increasing the temperature, we find a monotonic increase in the density of effectively free electrons, leading to a corresponding decrease of $\widetilde{r}_{s}$ and an increase in $\widetilde{Z}$. The observed monotonic increase of $\widetilde{\Theta}$ is less trivial. On the one hand, we have $\widetilde{\Theta}\sim\widetilde{r}_{s}^{2}$ 81, which, by itself, would indicate a decrease of $\widetilde{\Theta}$ with $T$. However, this effect is overridden by the relation $\widetilde{\Theta}\sim T$ in practice.
Table 1 clearly shows that the effectively free electrons in the interstitial region of isochorically heated copper are in the WDM regime 32. In nature, WDM occurs in a variety of astrophysical objects such as giant planet interiors 3, brown dwarfs 36, and white dwarf atmospheres 35. In the laboratory, these conditions are encountered on the compression path of a fuel capsule and its ablator in inertial confinement fusion experiments 4, 38, 6, and, in addition, they are relevant for material science, synthesis, and discovery 82, 83. Despite its fundamental importance for a gamut of applications, the rigorous theoretical description of WDM remains challenging: it must cover the complex interplay between effects such as partial ionization, Coulomb coupling, and quantum degeneracy and diffraction, which is notoriously difficult in practice 32, 33, 34. Our new simulation results thus imply that future XRTS experiments with isochorically heated copper constitute a suitable and highly controlled testbed for the benchmarking of theoretical methods and simulations.
@@ 4 Conclusions
Motivated by experimental capabilities at modern FEL facilities, we have performed a detailed study of the changes in the electronic DSF in fcc copper due to isochoric heating. Our new LR-TDDFT simulations show that the heating induces a prominent plasmon feature that eventually becomes dominant over the $d$-band-signal. This is in marked contrast to the ground state, where the plasmon is strongly damped by the presence of $d$-band excitations, and where it does not exhibit a meaningful dispersion with respect to the wavenumber $q$. Indeed, we have shown that at $T\geq 4~{}{\rm eV}$, the plasmon dispersion follows the familiar Bohm-Gross type relation, and the plasma frequency $\omega_{p}$ substantially increases with the temperature. This has been explained by the accumulation of effectively free electrons in the interstitial region between the ions. This, in turn, is a consequence of the availability of the electrons in the $d$-orbitals for filling states in the quasicontinuum when the temperature is increased. Interestingly, the reported behavior of the plasmon in copper is fundamentally different from isochorically heated aluminium (Al), where LR-TDDFT simulations have shown that heating to a few electronvolts (at $T\lesssim 7~{}{\rm eV}$) causes a red shift of $\omega_{p}$ 25. In this regard, we note that Al has 3 valence electrons from 3p3s orbitals and further ionization of the electrons from 2p requires temperatures about $70~{}{\rm eV}$. Therefore, in contrast to copper, thermal excitations in Al are not accompanied by an increase in the density of valence electrons with the increase in the temperature at $T\lesssim 10~{}{\rm eV}$.
@@
In this regard, we propose to use future XRTS experiments with isochorically heated copper as a controlled testbed for the rigorous assessment of different theoretical models and simulation tools for the description of WDM. Ideally, one might infer the temperature of the heated sample based on the model-free imaginary-time thermometry approach 11, 12 as a first step. Indeed, even the inference of comparably moderate temperatures of $T\sim 1-10\,$eV is expected to be feasible using the new high-resolution set-up that has recently been demonstrated at the European XFEL in Germany 24. Second, we propose to carry out measurements at multiple scattering angles (i.e., multiple wavenumbers $q$) to observe the plasmon dispersion $\omega(q)$ and, in this way, to infer the effective charge state $\widetilde{Z}$ and the effective density parameter $\widetilde{r}_{s}$. Since the ambient density is a-priori known, this will give one full access to the most relevant parameters of the system. This is an important advantage over shock experiments, or XRTS measurements with backlighter sources 39, where one or multiple of these parameters can only be inferred on the basis of the theoretical models which we aim to test in the first place. Isochoric heating can be achieved by employing a 400 nm optical short-pulse laser on thin targets and combining it with the delayed FEL probe 84, 85. Alternatively, one can use X-ray pump and X-ray probe beams that are separated in both colour and time. Proper time separation allows heating to be completed by the pump before the target is probed. Such X-ray pump X-ray probe experiments can be performed at European XFEL 86 and at SACLA in Japan 87, 88. A detailed discussion of the feasibility of measuring thermal excitations in the DSF of isochorically heated targets is provided in Ref. 25.
Such a hypothetical XRTS dataset can then be used to benchmark the zoo of available theoretical methods such as the widely used effective chemical models 19, 89, 90. A particularly interesting question is the assessment of XC functionals in KS-DFT simulations, including the resolution of explicitly thermal XC effects 91, 92, 93, 94, 95, 96. Moreover, frequency-resolved inelastic X-ray scattering data for a set of finite wavenumbers $q$ will be ideally suited to gauge the accuracy of different XC-kernels in LR-TDDFT calculations, including a rigorous assessment of the popular adiabatic approximation 40, 44, 97. As a WDM testbed, the observed isotropy of the DSF of electrons in copper is advantageous for achieving unambiguous conclusions in the assessment of various models.